Fermi National Laboratory
 

Operation Summary for May 18, 2011 - May 20, 2011

 

Wednesday May 18

The day shift began with Operations monitoring the Tevatron (TEV) store 8746 (49.92E30) and the Recycler stash (423.6E10), with the Antiproton Source (Pbar) stacking (16.9E10), with the Main Injector (MI), Linac, and Booster operating normally, with NuMI off to replace its target, with MiniBooNE off for access, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, MiniBooNE came out of access at 11:27 AM. Operations established beam to MiniBooNE at 11:45 AM. TEV personnel completed their end-of-store (EOS) crystal collimator studies at 12:10 PM. Operations terminated store 8746 at 12:57 PM, put the TEV into shot setup, and established store 8747 at 2:14 PM with an average initial luminosity (AIL) of 397.25E30. MI personnel conducted g-2 studies during the shift.

During the evening shift, per instructions, Operations lowered MiniBooNE’s beam intensity to around 3 to 4E16 protons per hour.

Thursday May 19

The midnight shift began with Operations monitoring store 8747 (111.62E30) and the stash (365.6E10), with Pbar stacking (22.2E10), and with MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8747 at 1:03 AM, put the TEV into shot setup, and established store 8748 at 2:24 AM with an AIL of 400.2E30.

During the day shift, TEV personnel began crystal collimator studies at 11 AM. Operators reported some short downtimes due to Linac RF stations (LRF1 and LRF2) trips and also to MI power supply (I:VT837) trips on ground fault indications.

During the evening shift, Operations reported that the crystal collimator studies concluded at 5:20 PM. Operations terminated store 8748 at 5:24 PM, put the TEV into shot setup, and established store 8749 at 6:46 PM with an AIL of 353.37E30. Operations reported some downtime when the Pbar autotune program lost communications with the antiproton BPMs; restarting the console resolved these problems.

Friday May 20

The midnight shift began with Operations monitoring store 8749 (156.16E30) and the stash (164.3E10), with Pbar stacking (19.9E10), and with MiniBooNE taking beam.

During the midnight shift, TEV personnel began EOS hollow beam studies at 6 AM. Operations reported minor downtime due to Booster RF trips. Operations reported a network storm at 7:04 AM that lasted about twenty minutes. Operations reported that the TEV FBI node went down several times and needed reboots.

The Plans for Friday and the Weekend

The plans for today are to terminate store 8749 after the TEV hollow beam study ends. The Run Coordinator will then allow CDF a three-hour access for necessary repairs and she will allow Booster studies. The rest of the day and the weekend will consist of stacking, stashing, and storing.
 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/20/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory